What are the x and y intercepts:
2x-3y=-10

Answer by nyc_function(2741)

To find x-intercept, set y = 0 and solve for x.

2x -3(0)= -10
2x = -10
x = -10/2
x = -5

To find y-intercept, set x = 0 and solve for y.
2(0)-3y = -10
-3y = -10

y = -10/-3

y = 10/3

The x-intercept is x = -5, which can also be written as (-5,0).
The y-intercept is y = 10/3, which can also be written as (0,10/3).